OpenStudy (moneybaby980):

The students at Midtown Middle school sold flowers as a fundraiser in September and October. In October, they charged $1.50 for each flower. The October price was a 20% increase of the September price. Part A What was the price of the flowers in September? Enter your answer in the box. $

OpenStudy (moneybaby980):

help

OpenStudy (phi):

call the unknown price x they say the price (i.e. x) plus 20% of x adds up to 1.50

OpenStudy (phi):

20% is the same as 20/100 or 0.2 as a decimal x + 0.2x = 1.50 or x * 1.2 = 1.50 I would divide both sides by 1.2 to find out what x is

OpenStudy (moneybaby980):

1.25

OpenStudy (moneybaby980):

is that right @phi

OpenStudy (phi):

yes. you can check by finding 20% of 1.25 0.2*1.25= 0.25 if we add this to the price: 1.25 + 0.25 = 1.50 we get the Oct price. looks ok